# Automated Statistical Testing and Certification of a Reliable Model-Coupling Server for Scientific Computing

### v1

@@ Abstract
Sequence-based specification and usage-driven statistical testing are designed for rigorous and cost-effective software development, offering a semi-formal approach to assessing the behavior of complex systems and interactions between various components. This approach is particularly valuable for scientific computing applications in which comprehensive tests are needed to prevent flawed results or conclusions. As scientific discovery becomes increasingly more complex, domain scientists couple multiple scientific computing models or simulations to solve intricate multiphysics and multiscale problems. These model-coupling applications use a hardwired coupling program or a flexible web service to link and combine different models. In this paper, we focus on the quality assurance of the more elastic web service via a combination of rigorous specification and testing methods. The application of statistical testing exposes problems ignored by pre-written unit tests and highlights areas in the code where failures might occur. We certify the model-coupling server controller with a derived reliability statistic, offering a quantitative measure to support a claim of its robustness.
@@ 4.3 Enumeration and Mealy Machine
With the completed enumeration, a Mealy machine is created using canonical sequences as states. Rows of the enumeration table define transitions among states and outputs on the arcs. The Mealy machine for the Data Exchange Controller, in Figure 2, is used to generate test cases and define the test oracle. A Markov-chain usage model is created by assigning probabilities to every arc from every state.
@@ 5.2 Results
Tests are run on three versions of the Data Exchange Controller. The old version is a prototype to demonstrate the idea of a flexible model-coupling server and was written without formal requirements. The new version is written with requirements derived from the old version, and a newer, fixed version is included with bugs identified and fixed during statistical testing. The bugs found during testing the new version are shown in Table 5, with fragments of the failed sequences shown in the first column.
@@
$C_{f}$ The implementation of the function associated with $C_{f}$ is missing information in the $var\_id$ field. Session was created regardless of the missing information. Session creation function did not check for missing information. New checks were added to the session creation function to check session creation request for flags and relevant parameters.

### v2

@@ Abstract
Sequence-based specification and usage-driven statistical testing are designed for rigorous and cost-effective software development, offering a semi-formal approach to assessing the behavior of complex systems and interactions between various components. While this approach has been successfully applied to a number of domains ranging from medical devices to scientific instrumentation, it is particularly valuable for scientific computing applications in which comprehensive tests are needed to prevent flawed results or conclusions. As scientific discovery becomes increasingly more complex, domain scientists couple multiple scientific computing models or simulations to solve intricate multiphysics and multiscale problems. These model-coupling applications use a hardwired coupling program or a flexible web service to link and combine different models. In this paper, we focus on the quality assurance of the more elastic web service by automatically generating, executing, and evaluating 5,204 test cases via a combination of rigorous specification and testing methods. The application of statistical testing exposes problems ignored by pre-written unit tests and highlights areas in the code where failures might occur. We certify the model-coupling server controller with a derived reliability statistic, offering a quantitative measure to support a claim of its robustness.
@@ 4.3 Enumeration and Mealy Machine
With the completed enumeration, a Mealy machine is created using canonical sequences as states. Rows of the enumeration table define transitions among states and outputs on the arcs. The Mealy machine for the Data Exchange Controller, in Figure 2, is used to generate test cases and define the test oracle. The Mealy machine is derived manually from the enumeration following a defined procedure [9, 15]. A Markov-chain usage model is created by assigning probabilities to every arc from every state.
@@ 5.2 Results
Tests are run on three versions of the Data Exchange Controller. The old version is a prototype to demonstrate the idea of a flexible model-coupling server and was written without formal requirements. The new version is written with requirements derived from the old version, and a newer, fixed version is included with bugs identified and fixed during statistical testing.
The bugs found during statistical testing of the new version are shown in Table 5, with fragments of the failed sequences shown in the first column. Reasons of failure were identified with human inspection.
@@
$C_{f}$ The function associated with $C_{f}$ is missing input validation. Session was created with missing information. Function did not check for missing information. Input validation added to the function to check request for valid parameters.